OpenStudy (anonymous):

Solve x+x/(x-2)=2/(x-2)

OpenStudy (zzr0ck3r):

\[x+\frac{x}{x-2}=\frac{2}{x-2}\]?

OpenStudy (zzr0ck3r):

multiply everything by x-2 (x-2)x+x=2 x^2-2x+x=2 x^2-x=2 x^2-x-2=0 x^2+x-2x-2 x(x+1)-2(x+1)=0 (x+1)(x-2)=0 x=-1 or x=2

OpenStudy (zzr0ck3r):

but x cant be 2

OpenStudy (zzr0ck3r):

so x = -1

OpenStudy (anonymous):

Dam your a life saver